SUB-DOPPLER SPECTROSCOPY OF MIXED STATE LEVELS IN CH_{2}

\underline{CHIH-HSUAN CHANG}, GREGORY E. HALL AND TREVOR J. SEARS, Chemistry Department, Brookhaven National Laboratory, Upton, New York, 11973.

Saturation dip spectroscopy has been used to measure rovibronic transitions in the b̃^{1}B_{1} ← ã^{1}A_{1} band system of CH_{2} at sub-Doppler resolution. The radical was made by 308 nm excimer laser photolysis of a slowly flowing, low-pressure, sample of ketene (CH_{2}CO), optionally with added inert buffer gas. Typical observed linewidths in the pure precursor are approximately 8 MHz (FWHM), due to a combination of collisional lifetime and pump-probe beam crossing angle. Due to the non-zero ^{1}H proton nuclear spin, CH_{2} exists as two distinct variants, ortho-CH_{2} with I_{H} = 1 and para-CH_{2} with I_{H} = 0. In ortho-CH_{2}, each rotational level consists of a triplet of hyperfine components corresponding to levels with F = J, J ± 1. Most singlet CH_{2} transitions show unresolved hyperfine structure in our experiment, since the largest splitting is due to I.J coupling, typically of the order of kHz. However, a small number of rotational levels in the v = 0 level of the lower ã state are known to be perturbed by accidentally near degenerate^{a} X̃^{3}B_{1}-state levels via spin-orbit coupling. Spectra involving such levels in ortho-CH_{2} exhibit resolvable triplet, I.S, hyperfine splittings, with the splittings providing a direct measure of triplet state character of the level. We have measured hyperfine splittings for a number of pairs of perturbed levels confirming and refining previous estimates of the singlet-triplet mixing coefficients. Measurements of the pressure-dependent saturation recovery rates with different collision partners can give new insights into dephasing, velocity-changing and inelastic collisions relevant to pressure broadening and intersystem crossing mechanisms. Acknowledgments: Work at Brookhaven National Laboratory was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy and supported by its Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences and Biosciences.

aU. Bley and F. Temps, J. Chem. Phys. 98, 1058-1072 (1993)